OpenStudy (anonymous):

\left( 3sqrt{10+4} \right)\left( 4sqrt{10-5} \right)

OpenStudy (anonymous):

\[\left( 3\sqrt{10+4} \right)\left( 4\sqrt{10-5} \right)\] simplify in radical expression.

myininaya (myininaya):

3*4=12 10+4=14 10-5=5 14(5)=70 \[12\sqrt{70}\]

OpenStudy (anonymous):

I tried that answer, not the right answer.

myininaya (myininaya):

it is the right answer

myininaya (myininaya):

are you sure you type the question in correctly?

OpenStudy (anonymous):

yes i am sure.

myininaya (myininaya):

you are entering this into like a online quiz?

OpenStudy (anonymous):

it has to simplified into radical form.

OpenStudy (anonymous):

No, it's online hw.

OpenStudy (anonymous):

It's the only one I can't get.

OpenStudy (anonymous):

Would you like me to show you the practice question? and give you the answer?

myininaya (myininaya):

you can if you want, but the above cannot be simplfied anymore since 70=2*5*7 there are no perfect squares in 70

OpenStudy (anonymous):

\[\left( 2\sqrt{6+3} \right) \left( 3\sqrt{6-4} \right)\]

myininaya (myininaya):

2(3)=6 6+3=9 6-4=2 6(3)sqrt(2) =18sqrt(2)

OpenStudy (anonymous):

answer is 24+\[24+\sqrt{6}\]

OpenStudy (anonymous):

Haha that's the answer it's giving me.

myininaya (myininaya):

24+sqrt(6)?

OpenStudy (anonymous):

Lol yea

myininaya (myininaya):

so this is multiplication right? not addition or subtraction or division? it is definitely multplication, right?

OpenStudy (anonymous):

multiplication for sure.

myininaya (myininaya):

ask your teachers. these solutions don't make sense

myininaya (myininaya):

there has to be a mistake or something

OpenStudy (anonymous):

It's annoying me. I always get 100% on the hw. Except for this one.

OpenStudy (anonymous):

It's due by midnight

myininaya (myininaya):

you have class tomorrow?

myininaya (myininaya):

does it show you how to work the problem, jason?

myininaya (myininaya):

if it does, we can try to follow the logic of the problem they did and see if we can get the answer they are looking for even though it will not be correct

OpenStudy (anonymous):

yea, but with weird examples that don't relate to it.

OpenStudy (anonymous):

Hold on, lemme get that for you.

OpenStudy (anonymous):

\[\left( \sqrt{5} +\sqrt{2}\right) \left( \sqrt{2} +3\right)\]

myininaya (myininaya):

sqrt(5)sqrt(2)+sqrt(5)*3+sqrt(5)sqrt(2)+sqrt(2)3 =sqrt(10)+3sqrt(5)+sqrt(10)+2sqrt(2) =2sqrt(10)+3sqrt(5)+2sqrt(2)

OpenStudy (anonymous):

answer is: \[\sqrt{10} +3\sqrt{5} + 2 + 3\sqrt{2}\]

myininaya (myininaya):

yes i made a mistke up there one sec

myininaya (myininaya):

sqrt(5)sqrt(2)+sqrt(5)3+sqrt(2)sqrt(2)+3sqrt(2) =sqrt(10)+3sqrt(5)+3sqrt(2)+2

myininaya (myininaya):

so maybe our problem was suppose to look like this

OpenStudy (anonymous):

What seems to be the problem then? :/

OpenStudy (anonymous):

On that practice question, I don't get how they got 24 +(sqrt)6

myininaya (myininaya):

ok lets pretend the problem was 3(sqrt(10)+4)2(sqrt(10)-5) 6(sqrt(10)+4)(sqrt(10)-5) 6(10-5sqrt(10)+4sqrt(10)-20) 60-sqrt(10)-20 40-sqrt(10 so this doesn't work

myininaya (myininaya):

so what did it say we are suppose to get?

OpenStudy (anonymous):

blank + (sqrt) blank

myininaya (myininaya):

ok let's look at that second one and see if we can get they got by making a problem sortof like but in the form of the example you gave

OpenStudy (anonymous):

the answer to that second one was 24+(sqrt)6

myininaya (myininaya):

omg i figured out the problem they meant

OpenStudy (anonymous):

....? :]

myininaya (myininaya):

so they have (2sqrt(6)+3)(3sqrt(6)-4)=24+sqrt(6)

OpenStudy (anonymous):

yea, how did you get that?

myininaya (myininaya):

okay try to entrt 100+srt(10)

OpenStudy (anonymous):

OMG!

OpenStudy (anonymous):

Thank you. Please elaborate?

myininaya (myininaya):

lol this is the problem they meant: (3sqrt(10)+4)(4sqrt(10)-5)

OpenStudy (anonymous):

ohhhhh my god. really? I didn't even see that. I'm sorry.

myininaya (myininaya):

how could you see it when it was written wrong or in the homework wrong

myininaya (myininaya):

you should totally tell your teacher of this mistake

OpenStudy (anonymous):

Thank you so much for the help, and I will do that. I have a quiz tmrw.

myininaya (myininaya):

do you understand how i got the answers? i don't expect you to understand how i was able to make the problem correct lol